OpenStudy (anonymous):

Variable y varies directly with variable x, and y = 45 when x = 9. What is value of y when x = 2? y = I will medal and fan for the best answer!!

OpenStudy (anonymous):

10 i do believe

OpenStudy (anonymous):

THank you I'll ask my teacher if that is correct

OpenStudy (anonymous):

my math isnt the best ever but thats what i would say

OpenStudy (anonymous):

was it correct?

OpenStudy (anonymous):

I don't know yet.. let me ask

OpenStudy (anonymous):

@samsmores

OpenStudy (anonymous):

i double checked im confident it is correct

OpenStudy (anonymous):

if 9*5=45(5x)then if x=2, 2*5=10.the answer is 10